# A template for the arxiv style

### v1

@@ A.2 Proof of Theorem 2
Lemma 3 (Adapted from Dasgupta and Gupta (2003), Lemma 2.2).
@@
Let $S\subset[D]$ with $S^{*}\not\subseteq S$ and set $U\vcentcolon=S\setminus S^{*}$. Then, under Assumptions 1, 3 and 4 we have the following properties of the SSR when regressing ${Y}$ onto ${X}_{S}$ in environments $v\in[E_{1}]$ and $w\in[E_{2}]$. Defining $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}$ it holds that $\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$. Here $r^{v}_{S}$ and $r^{w}_{S}$ are the residuals defined in Algorithm 3.
@@
where $r^{v}_{S}=\beta^{1}_{U}{X^{v}_{U}}+\varepsilon^{e}$. Because of the normality and independence assumptions, we find that $r^{v}_{S}$ follows a normal distribution with variance $\rho^{v}=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$. Furthermore $r^{v}_{S}$ can be assumed to be of zero mean, due to our inclusion of a column of constant ones in $X^{v}_{S}$. This implies that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(k)$ and we can define the equivalent expression for $w\in[E_{2}]$ by setting $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(k)$. ∎
@@
We continue with analyzing the quantity $\mathbf{P}(T(\boldsymbol{X},\boldsymbol{Y})>{Z_{\text{min}}}/{Z_{\text{max}}})$ and in particular try to understand the distribution of $T(\boldsymbol{X},\boldsymbol{Y})$ under Assumption 4. By Lemma 4 we know that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ for $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}$. This allows us to write:
@@
A.3 Proof of Theorem 3
First, we state two useful lemmas needed for the proof. We do not claim originality on those statements as those type of derivations may be found in literature on extremal events such as Embrechts et al. (2013). As we could, however, not find the precise statements needed, we prove them now.
@@ A.3 Proof of Theorem 3
We prove the lemma by showing that for any $c>1$ we have $\lim\limits_{E\to\infty}\mathbf{P}({W_{E}}/{Q_{E}}>c)=0$ and for any $c<1$ that $\lim\limits_{E\to\infty}\mathbf{P}({W_{E}}/{Q_{E}}<c)=0$. The statement of the lemma then follows from a union bound over the events $\{{W_{E}}/{Q_{E}}\not\in[1-\frac{1}{m},1+\frac{1}{m}\}_{m\in\mathbb{N}}$. Hashorva et al. (2012) show that there exists a series $b_{E}$ such that for $E\to\infty$ we have $b_{E}\to\infty$ and $(W_{E}-b_{E})$ converges to a distribution with support on $\mathbb{R}$. With that in hand we start by showing the case for $c>1$. First, choose $\delta>0$ such that $c>\frac{1+\delta}{1-\delta}$. With this we have that
@@
To identify when our method accepts $S$ as a set of potential causal parents, we have to understand the probability $\mathbf{P}(T_{S}>{\min_{e\in[E]}Z^{e}_{S}}/{\max_{e\in[E]}Z^{e}_{S}}$), which is the key quantity for the hypothesis test $\phi_{S}$, defined in Equation (9), that our Algorithm 3 uses. For notational convenience we set $Z_{\text{min}}\vcentcolon=\min\limits_{e\in[E]}Z^{e}$ and $Z_{\text{max}}\vcentcolon=\max\limits_{e\in[E]}Z^{e}$. For this proof we assume without loss of generality that $\rho^{v}>\rho^{w}$. We are investigating the behavior of $\mathbf{P}(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>{Z_{\text{min}}}/{Z_{\text{max}}})$ for $E\to\infty$, the statements of the theorem will then follow by Markov’s Inequality. By Lemma 4 we know that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ for $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{v}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{w}_{u})^{2}+(\sigma_{Y})^{2}$. This allows us to rewrite $T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})$ as:
@@ Appendix B Data Generation and Additional Results
In this section we describe the precise data generation mechanisms used in our experiments. Unless otherwise stated we fixed $E=30$, $D=6$ and $|S^{*}|=2$. In Section 6.1 and 6.2 we also experiment with data that was generated with a linear structural equation model (SEM). The SEM is defined by Equations (30)-(36) further below, with $\mathcal{D}(\sigma)$ being a distribution with zero mean and standard deviation $\sigma$. The specific choices of $\sigma$ for the individual experiments are specified in the following subsections. The graphical representation of the SEM is shown in Figure 9.
$X_{1}=\mathcal{D}(0,\sigma_{1})$ (30)
$X_{2}=X_{1}+\mathcal{D}(0,\sigma_{2})$ (31)
$X_{3}=0.3X_{1}+\mathcal{D}(0,\sigma_{3})$ (32)
$X_{4}=0.2X_{3}+\mathcal{D}(0,\sigma_{3})$ (33)
$Y=\beta_{2}X_{2}+\beta_{3}X_{3}+\mathcal{D}(0,\sigma_{Y})$ (34)
$X_{5}=0.1X_{2}+Y+\mathcal{D}(0,\sigma_{5})$ (35)
$X_{6}=Y+\mathcal{D}(\sigma_{6})$ (36)
B.1 Data from Section 6.1
@@
The data for a single run is generated in the following way. For each environment $e\in[E]$ we sample a vector of standard deviations $(\sigma^{e}_{1},\cdots,\sigma^{e}_{6})$, such that each entry is independently sampled from the uniform distribution on $[1,5]$. The support entries of $\beta^{e}$, which are given by $S^{*}=\{2,3\}$, are sampled in the same way. Similarly, a vector of means $(m_{1}^{e},\cdots,m_{6}^{e})$ is entry-wise sampled independently from the uniform distribution on $[-1,1]$. The entries of ${X}^{e}_{d}$ are then an i.i.d $n$-sample of $\mathcal{D}(m^{e}_{d},\sigma^{e}_{d})$, where the specific distribution is specified in the experiments, i.e., $\mathcal{D}(m^{e}_{d},\sigma^{e}_{d})$ is either a normal, uniform or student-t distribution with mean $m^{e}_{d}$ and standard deviation $\sigma^{e}_{d}$. The target is defined as ${Y}^{e}=\beta^{e}{X}^{e}+{\varepsilon}^{e}$, such that ${\varepsilon}^{e}$ is an i.i.d $n$-sample from $\mathcal{D}(0,2)$. The data of SEM is created similarly, but with the additional relations defined through Equation (30)-(36).
B.2 Data from Section 6.2
[…]
@@
B.3 Full Results of the Network Detection
Here we report the full counts of the experiments of Section 6.3. More precisely, let $\tilde{S}_{r,j}$ be the set of causal parents that LoLICaP with given sample size $n$ reported in run $r$ for target covariate $j$, then we define $M^{n}_{i,j}:=\sum_{r=1}^{500}\mathbf{1}\left\{i\in\tilde{S}_{r,j}\right\}$. The results of the experiments from Section 6.3 are then given by the following two matrices
$M^{20}=\begin{bmatrix}494&23&7&6&9&3\\ 43&489&130&13&19&17\\ 54&108&442&371&44&24\\ 109&90&120&362&361&42\\ 82&81&53&405&405&41\\ 4&8&3&8&6&490\end{bmatrix},M^{25}=\begin{bmatrix}498&56&3&3&3&3\\ 68&470&238&28&27&17\\ 87&193&337&320&97&43\\ 110&93&108&189&189&44\\ 96&72&58&227&227&38\\ 2&5&3&6&3&496\end{bmatrix}.$

### v2

@@ A.4 Proof of Theorem 2
Lemma 3 (Adapted from Dasgupta and Gupta [2003], Lemma 2.2).
@@
Let $S\subset[D]$ with $S^{*}\not\subseteq S$ and set $U\vcentcolon=S\setminus S^{*}$. Then, under Assumptions 1, 3 and 4 we have the following properties of the SSR when regressing ${Y}$ onto ${X}_{S}$ in environments $v\in[E_{1}]$ and $w\in[E_{2}]$. Defining $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{1}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{1}_{u})^{2}+(\sigma_{Y})^{2}$ it holds that $\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$. Here $r^{v}_{S}$ and $r^{w}_{S}$ are the residuals defined in Algorithm 3.
@@
where $r^{v}_{S}=\beta^{1}_{U}{X^{v}_{U}}+\varepsilon^{e}$. Because of the normality and independence assumptions, we find that $r^{v}_{S}$ follows a normal distribution with variance $\rho^{v}=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{1}_{u})^{2}+(\sigma_{Y})^{2}$. Furthermore $r^{v}_{S}$ can be assumed to be of zero mean, due to our inclusion of a column of constant ones in $X^{v}_{S}$. This implies that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(k)$ and we can define the equivalent expression for $w\in[E_{2}]$ by setting $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(k)$. ∎
@@
We continue with analyzing the quantity $\mathbf{P}(T(\boldsymbol{X},\boldsymbol{Y})>{Z_{\text{min}}}/{Z_{\text{max}}})$ and in particular try to understand the distribution of $T(\boldsymbol{X},\boldsymbol{Y})$ under Assumption 4. By Lemma 4 we know that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ for $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{1}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{1}_{u})^{2}+(\sigma_{Y})^{2}$. This allows us to write:
@@
A.5 Proof of Theorem 3
First, we state two useful lemmas needed for the proof. We do not claim originality on those statements as those type of derivations may be found in literature on extremal events such as Embrechts et al. [2013]. As we could, however, not find the precise statements needed, we prove them now.
@@ A.5 Proof of Theorem 3
We prove the lemma by showing that for any $c>1$ we have $\lim\limits_{E\to\infty}\mathbf{P}({W_{E}}/{Q_{E}}>c)=0$ and for any $c<1$ that $\lim\limits_{E\to\infty}\mathbf{P}({W_{E}}/{Q_{E}}<c)=0$. The statement of the lemma then follows from a union bound over the events $\{{W_{E}}/{Q_{E}}\not\in[1-\frac{1}{m},1+\frac{1}{m}\}_{m\in\mathbb{N}}$. Hashorva et al. [2012] show that there exists a series $b_{E}$ such that for $E\to\infty$ we have $b_{E}\to\infty$ and $(W_{E}-b_{E})$ converges to a distribution with support on $\mathbb{R}$. With that in hand we start by showing the case for $c>1$. First, choose $\delta>0$ such that $c>\frac{1+\delta}{1-\delta}$. With this we have that
@@
To identify when our method accepts $S$ as a set of potential causal parents, we have to understand the probability $\mathbf{P}(T_{S}>{\min_{e\in[E]}Z^{e}_{S}}/{\max_{e\in[E]}Z^{e}_{S}}$), which is the key quantity for the hypothesis test $\phi_{S}$, defined in Equation (5), that our Algorithm 3 uses. For notational convenience we set $Z_{\text{min}}\vcentcolon=\min\limits_{e\in[E]}Z^{e}$ and $Z_{\text{max}}\vcentcolon=\max\limits_{e\in[E]}Z^{e}$. For this proof we assume without loss of generality that $\rho^{v}>\rho^{w}$. We are investigating the behavior of $\mathbf{P}(T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})>{Z_{\text{min}}}/{Z_{\text{max}}})$ for $E\to\infty$, the statements of the theorem will then follow by Markov’s Inequality. By Lemma 4 we know that $P^{v}\vcentcolon=\frac{1}{\rho^{v}}\|r^{v}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ and $Q^{w}\vcentcolon=\frac{1}{\rho^{w}}\|r^{w}_{S}\|_{2}\sim\chi^{2}(n-|S|)$ for $\rho^{v}\vcentcolon=\sum_{u\in U}(\beta^{1}_{u})^{2}(\sigma^{1}_{u})^{2}+(\sigma_{Y})^{2}$ and $\rho^{w}\vcentcolon=\sum_{u\in U}(\beta^{2}_{u})^{2}(\sigma^{2}_{u})^{2}+(\sigma_{Y})^{2}$. This allows us to rewrite $T(\boldsymbol{X}_{E},\boldsymbol{Y}_{E})$ as:
@@ Appendix B Data Generation and Additional Results
In this section we describe the precise data generation mechanisms used in our experiments from Section 6. Unless otherwise stated we fixed $E=30$, $D=6$ and $|S^{*}|=2$. The data is generated from a linear structural equation model with different noise distributions, given by Equations (26)-(32) further below. Here $\mathcal{D}(\sigma)$ is a distribution with standard deviation $\sigma$ and zero mean. The specific choices of $\sigma$ for the individual experiments are specified in the following subsections. The graphical representation of the SEM is shown in Figure 4.
$X_{1}=\mathcal{D}(\sigma_{1})$ (26)
$X_{2}=X_{1}+\mathcal{D}(\sigma_{2})$ (27)
$X_{3}=0.3X_{1}+\mathcal{D}(\sigma_{3})$ (28)
$X_{4}=0.2X_{3}+\mathcal{D}(\sigma_{3})$ (29)
$Y=\beta_{2}X_{2}+\beta_{3}X_{3}+\mathcal{D}(\sigma_{Y})$ (30)
$X_{5}=0.1X_{2}+0.3Y+\mathcal{D}(\sigma_{5})$ (31)
$X_{6}=0.5Y+\mathcal{D}(\sigma_{6})$ (32)
Effects of Non-Normal Noise.
@@
The data for a single run within the noise misspecification experiment is generated in the following way. For each environment $e\in[E]$ we sample a vector of standard deviations $(\sigma^{e}_{1},\cdots,\sigma^{e}_{6})$, such that each entry is independently sampled from the uniform distribution on $[1,5]$. The support entries of $\beta^{e}$, which are given by $S^{*}=\{2,3\}$, are sampled in the same way. The data is then created with the additional relations defined through Equation (26)-(32). The noise distributions $\mathcal{D}(\sigma_{1}),\cdots,\mathcal{D}(\sigma_{6})$ for the covariates are Gaussian with standard deviations as described above, while the target noise distribution $\mathcal{D}(\sigma_{Y})$ is either a uniform, Student-t or Gaussian distribution (as indicated in the figures) with $\sigma_{Y}=1.1$.
Comparison with ICP
As L-ICP and ICP are similar in many regards, we chose for three simple experiments that highlights the differences. We set $\mathcal{N}(m,s)$ to be the standard normal distribution with mean $m$ and standard deviation $s$. In all three experiments we independently sample in $E=100$ environments $n=7$ observations $X^{e}_{1,i}\sim\mathcal{N}(0,\sigma)$, $X^{e}_{2,i}\sim\mathcal{N}(0,\sigma)$ and $Y^{e}_{i}=\beta^{e}X^{e}_{1,i}+\varepsilon^{e}_{i}$ with $\varepsilon^{e}_{i}\sim\mathcal{N}(0,1)$.
In the dense setting we set $\beta^{e}=1$ and sample in each environment $\sigma$ from a uniform distribution on $[1,5]$.
In the sparse setting we set $\beta^{e}=1$, as well as $s=1$ for $99$ out of the $100$ environments. In the last environment we set $s=3$.
In the ICP violation setting we follow the dense setting, but additionally sample $\beta^{e}$ independently from a uniform distribution on $[1,5]$.
Comparison with LiNGAM.
The data for a single run within the comparison to LiNGAM experiment is generated in the following way. For each environment $e\in[E]=[30]$ with $e\leq 15$ we set $(\sigma^{e}_{1},\cdots,\sigma^{e}_{6})=(2,\cdots,2)$ and the support entries of $\beta^{e}$, which are given by $S^{*}=\{2,3\}$, are set to $(1,1)$. For $e>15$ we set $(\sigma^{e}_{1},\cdots,\sigma^{e}_{6})=(c,\cdots,c)$ and $\beta^{e}_{2}=\beta^{e}_{3}=c$ for $c\in\{1,1.2,1.4,1.5,1.7,1.8\}$ in the uniform noise case, $c\in\{1,1.2,1.4,1.5,1.6\}$ for the Gaussian noise case and $c\in\{1,2,5,8\}$ for the scaled Student-t noise. To obtain the heterogeneity parameter $h$ we first apply Equation 7 to obtain $I_{\{2\}}=I_{\{3\}}$ (the relevant quantities to avoid false negatives) and then set $h=(I_{\{2\}})^{\frac{1}{4}}e^{1-(I_{\{2\}})^{\frac{1}{4}}}$ as also explained in Section 5.1. In this experiment we chose to not randomly sample the heterogeneity for a better control over it. The data is then created with the additional relations defined through Equation (26)-(32). The noise distributions $\mathcal{D}(\sigma_{1}),\cdots,\mathcal{D}(\sigma_{6})$ and also $\mathcal{D}(\sigma_{Y})$ are all uniform distributions for the uniform noise experiment, Gaussian distributions for the Gaussian noise experiment and scaled Student-t distributions for the last experiment. In all cases we set $\sigma_{Y}=1$. Note that the experiment with the scaled Student-t distribution is found in Appendix B.1.
@@
B.1 Additional Experiments
Adjusting the calibration under noise misspecification.
In Figures 1(a) and 1(b) of Section 6 we observed that under Student-t distributed noise, both the false negative and the false positive rate is adversely affected, in particular for larger sample sizes. In the following we show that this is not an inherent problem of our test statistic, but just due to the wrong calibration that assumes normal noise. If we adjust the target calibration $\alpha$ we can indeed recover a good performance as shown in Figure 5. How such an adjustment may be done in practice is unclear, and for that reason an important extension of L-ICP will be to find ways to calibrate the method without the normality assumption.
Figure 5: Under Student-t distributed noise, L-ICP achieves not the target calibration and this affects, in particular for larger samples, the performance. A near-optimal performance can be recovered if we adjust $\alpha$.
Comparison with LiNGAM under Student-t noise.
To further study the effect of noise misspecification on L-ICPs performance we perform an additional comparison with LiNGAM when the noise comes from a scaled Student-t distribution. The results are shown in Figures 6(a) and 6(b). We notice that if we set the degree of freedom to $3$, so under a strong noise misspecification, L-ICP cannot recover good performance, also with strong heterogeneity of the environments. For $10$ degrees of freedom, resulting in a weaker noise misspecification, L-ICP is again able to recover a good performance when there is sufficient heterogeneity in the environments.
(a) The results when the scaled Student-t distribution has $3$ degrees of freedom. While LiNGAM does make use of the strong non-normality, L-ICP can under the strong misspecifcation not recover a good performance, even when a lot of heterogeneity is present.
(b) The results when the scaled Student-t distribution has $10$ degrees of freedom. As the noise misspecification is less strong, L-ICP can recover again good performance with the presence of heterogeneity.
B.2 Full Results of the Network Detection
The dynamical system is formally defined by the following set of equations, where the superscript $t$ indicates a time index.
$X^{t+1}_{1}=0.9X_{1}^{t}+0.1X^{t}_{2}+\varepsilon^{t}_{1}$ (33)
$X^{t+1}_{5}=0.02X^{t}_{1}X^{t}_{4}+0.96X^{t}_{5}+\varepsilon^{t}_{5}$ (37)
$X^{t+1}_{6}=X^{t}_{6}+\varepsilon^{t}_{6}$ (38)
Here we report the full counts of the experiments of Section 6.1, additionally also when we use $n=20$ samples for L-ICP. More precisely, let $\tilde{S}_{r,j}$ be the set of causal parents that L-ICP with given sample size $n$ reported in run $r$ for target covariate $j$, then we define $M^{n}_{i,j}:=\sum_{r=1}^{500}\mathbf{1}\left\{i\in\tilde{S}_{r,j}\right\}$. The results of the experiments from Section 6.1 are then given by the following two matrices
$M^{20}=\begin{bmatrix}494&43&54&109&82&4\\ 23&489&108&90&81&8\\ 7&130&442&120&53&3\\ 6&13&371&362&405&8\\ 9&19&44&361&405&6\\ 3&17&24&42&41&490\end{bmatrix},M^{25}=\begin{bmatrix}498&68&87&110&96&2\\ 56&470&193&93&72&5\\ 3&238&337&108&58&3\\ 3&28&320&189&189&6\\ 3&27&97&227&227&3\\ 3&17&43&44&38&496\end{bmatrix}.$
As PCMCI does not naturally group the environments, we run PCMCI over $30$ individual intervals of length $n=25$ in each of the $500$ runs. The complete counts for PCMCI are:
$N^{25}=\begin{bmatrix}10778&1906&2209&1996&2033&1868\\ 2422&12439&2035&1835&1830&1870\\ 1766&2095&12816&2239&1852&1809\\ 1875&1816&2201&12874&3273&1838\\ 1909&1848&2000&3030&12980&1960\\ 1857&1832&1851&1898&1879&9968\end{bmatrix}$
Based on the ground truth graph we picked a threshold of $1994$ and report the edge from $i$ to $j$ if $N^{25}_{i,j}>1994$.